OpenStudy (anonymous):

use the method of substitution to solve the system of linear equations. 7a=3-b 7a=4-b

OpenStudy (anonymous):

No solutions... think about it. If \(7a=3-b\) and \(7a=4-b\), since we know \(7a=7a\), that means \(3-b=4-b\). Now, if you add \(b\) to both sides, you end up with \(3=4\) -- which isn't true.